[Hallc_running] [New Logentry] RC Daily Update

Douglas Higinbotham doug at jlab.org
Fri Sep 16 10:22:38 EDT 2022

Due to some poor planning by the RF group, the key people aren’t here next week, so the new plan (as of a few minutes ago after texting with Camille) is they are going to push for full current swing shift today!

Thus Carlos please try to get the hall locked up as soon as possible.   (i.e. be sure the Hall is locked up by swing shift if at all possible)

And second, please let accelerator do there tuning and RF as long as they needed this evening.
(i.e. don’t bug them putting in targets, etc. until they finish getting the RF and high current beam setup)

Best Wishes,

From: Hallc_running <hallc_running-bounces at jlab.org> on behalf of cyero at jlab.org <cyero at jlab.org>
Date: Friday, September 16, 2022 at 10:10 AM
To: hallc_running at jlab.org <hallc_running at jlab.org>, weinstei at jlab.org <weinstei at jlab.org>
Subject: [Hallc_running] [New Logentry] RC Daily Update
RC Daily Update<https://logbooks.jlab.org/entry/4042342>

  *   Edit<https://logbooks.jlab.org/entry/4042342/edit?destination=email/send>
  *   Delete<https://logbooks.jlab.org/entry/4042342/delete?destination=email/send>

Lognumber 4042342<https://logbooks.jlab.org/entry/4042342>. Submitted by cyero<https://logbooks.jlab.org/user/cyero> on Fri, 09/16/2022 - 10:07<https://logbooks.jlab.org/entries?start_date=1663333676&end_date=1663340876&book=HCLOG>.
Entry Makers:
Carlos Yero

Hall C Status:
1) Target cool down in progress by D. Meekins

2) HMS/SHMS Rotation certification in progress by S. Lassiter
(will set spectrometers to CaFe initial settings after certification completed)

3) Hall C Fast Raster to be restored by B. Gunning at 10 AM today

4) Hall C Walkthrough by J. Beaufait to be done from ~ 2 - 3 pm today

5) If the above points go smoothly, Hall C to be swept this afternoon

Here is the detailed CaFe Run Plan: https://hallcweb.jlab.org/wiki/images/5/52/Cafe_runplan_10_6GeV_original...<https://hallcweb.jlab.org/wiki/images/5/52/Cafe_runplan_10_6GeV_originalOptics_official.pdf>

Accelerator Status:
1) Will try to re-establish beam to Hall C starting today

2) Will try to ramp-up to max current (80 uA) on Monday

No need for RC meeting today, unless requested.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220916/d2c98384/attachment-0001.html>

More information about the Hallc_running mailing list